Physical Chemistry

The number of unit cell in 10 gram of NaCl. ..................................................................................

Manu Tyagi
8 Years agoGrade

1 Answer

Arun
8 Years ago
Dear Manu
 
Each unit cell is made of 4 Na+ and 4 Cl- ? ions.
This means, each unit cell is made of 4 formula units.
No.of moles of  NaCl(n) = mass/molecular mass of NaCl
?10/58.5=0.17 moles
1 mole = 6.022 * 1023 formula units
0.17 moles = 6.022 * 1023 * 0.17 formula units, which is equal to 1.02 * 10-8formula units.
1 unit cell has 4 formula units of NaCl.
4 formula units = 1 unit cell
1.02 * 10-8 formula units = 2.55 * 10-9 unit cells.
Regards
Arun (askIITians forum expert)